OpenStudy (anonymous):

easiest way to solve this 4got how to solve use quadratic formula to solve this question round solutions to nearest hundereth if neccesary 1) -3x^2-4x+10=0 2) x^2+6x-100=0

OpenStudy (anonymous):

\[x^2+6x=100\] \[(x+3)^2=100+9=109\] \[x+3=\pm\sqrt{109}\] \[x=-3\pm\sqrt{109}\]

OpenStudy (anonymous):

THANK YOU Do you know the first one

OpenStudy (anonymous):

\[ -3x^2-4x+10=0\] \[3x^2+4x-10=0\] \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] witih \(a=3,b=4,c=-10\)